OpenStudy (anonymous):

The width of a rectangle is 3 feet, and the diagonal is 5 feet. What is the area of the rectangle?

OpenStudy (anonymous):

a^2 + b^2= c^2 (^2 means squared) b^2 = 5^2 - 3^2 b^2 = 25-9 =16 b = 4 width = 4 feet Area=l x w = 3 x 4 =12 square feet